shift sum 6/11/99 eve (one more attempt)

Theodoros Alexopoulos (theoalex@fnal.gov)
Sat, 12 Jun 1999 11:06:49 -0500 (CDT)

Shift Summary for evening shift, Jun. 11, 1999

On Shift: AE, TA
Also EB

Hours of beam delivered at all intensities: 3.5h
Hours of beam delivered at nominal intensity: 3.0h
Hours of E832 data (special runs not included): 3,5h
Hours of special runs (laser scans, muon runs...): 1.0h

Accounting of lost time: None

========================================================================
Runs ended this shift:

Run 13539 - Laser scan
Run 13541 - 59K events - BEAM_832
Run 13542 - 2.7M events - BEAM_832 (Run in progress)
This run looks great, it's the first real good BEAM_832 run!
Average SEM=1.5E12/spill
Here is a sample of the yields:
Qtty ON-LINE OFF-LINE OFF/ON ratio
------------------------------------------------------------------------
VAC(+-): 39743.55 +- 318.81 17751.67 +- 136.47 0.447 +-0.005
REG(+-): 69825.30 +- 279.78 28739.15 +- 170.41 0.412 +-0.003
VAC(00): 9014.56 +- 174.65 3831.08 +- 66.01 0.425 +-0.011
REG(00): 13629.16 +- 129.22 6385.12 +- 80.24 0.468 +-0.007

Tapes brought to FCC this shift: None

========================================================================
Accesses during shift: None

========================================================================
Overview of shift:
Laser scan and E832 datataking

========================================================================
Description of Shift:

16:00 Because thunderstorm triggered interlocks all over
the site, operators wait until storm is away before
they try to reset things.

17:45 TeV is off due to Cryo problems.

run spill time comment
------------------------------------------------------------------------
13539 1 19:00 start long laser scan
19:30 MCR called: beam shortly

Ed compars yields between runs
9247 (KTeV-data) vs 13524 (KTeV99-data):
Finds: charged mode yield increases by 44% !!!
neutral mode yield increases by 14% !
But VAC/REG is consistent for each mode.
13541 1 20:30 beam is on. sdaq gives a NM3S ramping alarm.
NM3S doesn;t ramp during the spill
Call the MCR. They run NM3S in a staedy state,
since they couldn't fix the ramping feature.
They are working on this problem.
21:10 ktev2 crashed. We unload the tapes (which
are more 1/2 filled). Do a fresh start.
13542 21:30 start a new run. SEM~1.6E12. Spill quality
looks much better in this run. Also the livetime
is now ~85%, compare to 75% at the same KTeV intensity.
23:30 sdaq ktev1a alarm. ktev1a is getting full 99%.
free space by removing some old files.

Looked at several monitoring kumacs. Many of them
need to be updated/corrected. Mail sent to
appropriate persons. People should go through their
kumacs and redefine the nominal ones.

========================================================================
Items to Follow up on Next Shift

Keep going with this BEAM_832 datataking.
Following first run stop after mifnight, do a muon run
and flip polarity of NM4AN for next hadron run.
Collect muons for 1 hour using BEAM_MUON_4E12.